# Worldsheet patching, 1-form symmetries, and “Landau-star” phase transitions

### v1

@@ VI.4 Application to experiments in quantum devices
Recently it was possible to realize the toric code ground state experimentally [52, 53]. A key question is how to efficiently detect deconfinement in experiments of this type by using measurements of the spins. One approach is the Fredenhagen-Marcu order parameter, which involves a ratio of correlation functions of “bare” Wilson line operators [24, 20, 22]. However this involves measuring exponentially small quantities (because of nonuniversal short-distance effects close to the Wilson line), implying a signal-to-noise problem in the large size limit. Therefore we suggest that an effective method may be instead to directly define emergent string correlators by applying the configuration “repair” idea of Ref. [14] to 2D quantum measurements, rather than 3D Monte Carlo data. Below we give a schematic indication.
@@ VII Outlook
Inside the deconfined phase, there are two emergent one-form symmetries, related by e–m duality. One of these symmetries survives at the Higgs transition, and the dual symmetry survives at the confinement transition. The present approach only gives access to one set of string operators at a time, because it is necessary to pick a basis (electric or magnetic) in order to formulate the partition function as a membrane ensemble. The phase diagram of the standard $\mathbb{Z}_{2}$ gauge theory has an exact self-duality, so it is straightforward to change between the two bases. For more general models (e.g. models in the spatial continuum) the analog of this basis change may be more nontrivial.
Finally let us discuss the multicritical point in Fig. 1. In addition to the confinement and Higgs transition lines, the phase diagram of the gauge theory contains a self-dual multicritical point (MCP) where they meet. The MCP is not yet understood as a conformal field theory, and is an interesting target for further study.
Since the MCP is the meeting point of two critical lines with Ising exponents, at first sight a natural guess would be that it has XY exponents [20]. More recently, this has been argued by Ref. [27], on the basis of the numerical similarity of exponents observed in Ref. [14]. However, challenges to this theoretical interpretation were discussed in Refs. [14, 28]. To begin with, the fixed point associated with the MCP cannot be the fixed point of the XY model (for example because the adjacent phases do not match). In addition, the MCP cannot be an orbifold [^12] Here orbifolding refers to gauging with a flat gauge field (which leaves correlators of gauge-invariant operators unchanged and eliminates non-gauge-invariant operators from the spectrum). The term is used in a different sense in some other contexts [56]. of the XY fixed point (i.e. an “$*{}^{*}$ transition”) as this would not give the correct universal properties of the topologically ordered phase immediately adjacent to the MCP [14].
Ref. [28] has directly demonstrated (using simulations) that there is no U(1) current operator at the MCP CFT, showing that the MCP does not have an emergent global U(1) symmetry. By contrast, the usual “$*{}^{*}$” transition has a U(1) symmetry.
@@ Acknowledgements.
We thank Snir Gazit, Lior Oppenheim, Zohar Ringel, and Dominic Williamson for discussions. A.S. acknowledges support by MCIN/AEI/10.13039/501100011033 through project PID2022-139191OB-C31 “ERDF A way of making Europe”.
@@ A.1 Lengthscales close to the self-dual MCP
This is the characteristic lengthscale for the crossover from the universality class of the MCP to universality class of the $*{}^{*}$ confinement transition. (Analogous statements apply for the crossover on the Higgs line.) $\ell(y)$ is also the lengthscale associated with patches, i.e. the scale beyond which larger patches become exponentially rare.

### v2

@@ VI.4 Application to experiments in quantum devices
Recently it was possible to realize the toric code ground state experimentally [52, 53]. A key question is how to efficiently detect deconfinement in experiments of this type by using measurements of the spins. One approach is the Fredenhagen-Marcu order parameter, which involves a ratio of correlation functions of “bare” Wilson line operators [24, 20, 22, 54]. However this involves measuring exponentially small quantities (because of nonuniversal short-distance effects close to the Wilson line), implying a signal-to-noise problem in the large size limit. Therefore we suggest that an effective method may be instead to directly define emergent string correlators by applying the configuration “repair” idea of Ref. [14] to 2D quantum measurements, rather than 3D Monte Carlo data. Below we give a schematic indication.
@@ VII Outlook & comments on the MCP
An alternative tool for diagnosing deconfinement is the Fredenhagen-Marcu order parameter [24, 22], whose critical behavior was recently studied numerically [54]. This observable uses a “bare” string operator, rather than a dressed one, and as a result generally suffers an exponential signal-to-noise problem in Monte Carlo. However the scaling of the FM order parameter could be understood via an extension of the arguments in Sec. V and the appendices, by writing the bare string operator as the product of the dressed operator and a contribution from patches.
In this paper we have considered the gauge theory with periodic boundary conditions, but it would also be interesting to explore boundary phenomena. Recently it has been argued that Higgs phases can often be thought of as symmetry-protected topological states protected by both conventional and higher-form symmetries [57], with associated boundary states. (See Refs. [58, 59] for other approaches to distinguishing Higgs and confining regimes.)
Inside the deconfined phase, there are two emergent one-form symmetries, related by e–m duality. One of these symmetries survives at the Higgs transition, and the dual symmetry survives at the confinement transition. The approach in the present paper only gives access to one set of string operators at a time, because it is necessary to pick a basis (electric or magnetic) in order to formulate the partition function as a membrane ensemble. The phase diagram of the standard $\mathbb{Z}_{2}$ gauge theory has an exact self-duality, so it is straightforward to change between the two bases. For more general models (e.g. models in the spatial continuum) the analog of this basis change may be more nontrivial.
VII.1 Comments on self-dual multicritical point
Finally let us discuss the multicritical point in Fig. 1. In addition to the confinement and Higgs transition lines, the phase diagram of the gauge theory contains the self-dual multicritical point (MCP) where they meet. The MCP is not yet understood as a conformal field theory, and is an interesting target for further study.
Since the MCP is the meeting point of two critical lines with Ising exponents, at first sight a natural guess would be that it has XY exponents [20]. More recently, this has been argued by Ref. [27], on the basis of the numerical similarity of exponents observed in Ref. [14]. However, challenges to this theoretical interpretation were discussed in Refs. [14, 28]. To begin with, the fixed point associated with the MCP cannot be the fixed point of the XY model (for example because the adjacent phases do not match). In addition, the MCP cannot be an orbifold [^13] Here orbifolding refers to gauging with a flat gauge field (which leaves correlators of gauge-invariant operators unchanged and eliminates non-gauge-invariant operators from the spectrum). The term is used in a different sense in some other contexts [60]. of the XY fixed point (i.e. an “XY transition”) as this would not give the correct universal properties of the topologically ordered phase immediately adjacent to the MCP [14].
Ref. [28] has directly demonstrated (using simulations) that there is no U(1) current operator at the MCP CFT, showing that the MCP does not have an emergent global U(1) symmetry. By contrast, the usual “XY” transition has a U(1) symmetry.
@@ VII.1 Comments on self-dual multicritical point
Instead, any continuum description needs to take account of the braiding statistics in the adjacent deconfined phase. The simplest framework that can account for the anyonic braiding statistics of $\mathbb{Z}_{2}$ topological order using continuum (as opposed to lattice) field theory is a mutual Chern-Simons (or “BF”) theory, with two U(1) gauge fields, $a^{1}$ and $a^{2}$, with a Chern-Simons term ${i\frac{k}{2\pi}\epsilon_{\mu\nu\lambda}a_{\mu}^{1}\partial_{\nu}^{\phantom{1}}a_{\lambda}^{2}}$ at ${k=2}$. These gauge fields are coupled to two matter fields, whose quanta represent e and m anyons respectively [61, 62, 63, 64]:
$\mathcal{L}_{\rm CS}=$ $\,\frac{i}{\pi}\epsilon_{\mu\nu\lambda}a^{1}_{\mu}\partial_{\nu}a^{2}_{\lambda}+\frac{1}{2}\sum_{I=1,2}|(\partial-ia^{I})z^{I}|^{2}$
$+$ $\frac{m^{2}}{2}\sum_{I=1,2}|z^{I}|^{2}+\lambda\left(|z^{1}|^{4}+|z^{2}|^{4}\right)+\lambda^{\prime}|z^{1}|^{2}|z^{2}|^{2}+\ldots$
However, if we want a putative description of the MCP, then we must also add monopole operators $\mathcal{M}^{(I)}$, for each of the gauge fields $I=1,2$, to this continuum Lagrangian. This is necessary to avoid an unphysical enlargement of symmetry, or equivalently to remove an unwanted degeneracy in the spectrum of anyons:
$\mathcal{L}=\mathcal{L}_{\rm CS}+\kappa\sum_{I=1,2}\left(\mathcal{M}^{(I)}+\mathcal{M}^{(I)*}\right).$ (22)
This issue was commented on in Ref. [14]. In App. C we review the Chern-Simons approach in more detail in order to clarify why the multicritical point is challenging to describe using continuum field theory. (See a very recent preprint for a discussion of $\mathbb{Z}_{k}$ for larger values of $k$ [65].) App. C also discusses a (speculative) alternative way to remove the unwanted anyon degeneracy, using a replica-like limit in the Chern-Simons theory.
@@ Acknowledgements.
We thank Snir Gazit, Lior Oppenheim, Zohar Ringel, Balt van Rees, Slava Rychkov, Chong Wang, and Dominic Williamson for discussions. A.S. acknowledges support by MCIN/AEI/10.13039/501100011033 through project PID2022-139191OB-C31 “ERDF A way of making Europe”.
@@ A.1 Lengthscales close to the self-dual MCP
This is the characteristic lengthscale for the crossover from the universality class of the MCP to universality class of the Ising confinement transition. (Analogous statements apply for the crossover on the Higgs line.) $\ell(y)$ is also the lengthscale associated with patches, i.e. the scale beyond which larger patches become exponentially rare.